OpenStudy (anonymous):

the line through (0,-2) and (4,6) in point-slope form. please show the work , thanks(: the line through (0,-2) and (4,6) in point-slope form. please show the work , thanks(: @Mathematics

OpenStudy (anonymous):

\[slope = \frac{y2-y1}{x2-x1} = \frac{6-(-2)}{4-0} = \frac{8}{4} = 2\] line equation: y=2x+b plug in one of the points you have (0,-2) or (4,6) to find b: -2 = 2(0) + b b = -2 Therefore, y = 2x - 2 is that right ?

jimthompson5910 (jim_thompson5910):

m=(y2-y1)/(x2-x1) m=(6-(-2))/(4-0) m=(8)/(4-0) m=(8)/(4) m=2 --------------------------------------------- y-y1=m(x-x1) y-(-2)=2(x-0) y+2=2(x-0)

OpenStudy (anonymous):

is y+2=2x the same thing?

jimthompson5910 (jim_thompson5910):

yes, but I figured I'd leave it in y - y1 = m(x-x1) form

OpenStudy (anonymous):

alright thanks (: